In (R, U) usual topology, find
A°,Z°, ext (A), ext (Z), b(A), b (Z)
Where A = {1,2} and Z set of all integer numbers
1
Expert's answer
2020-11-13T14:59:58-0500

Consider the topological space "(\mathbb R, U)" with usual topology "U". Since there is no basic neigbourhood "(1-\epsilon, 1+\epsilon)" such that "(1-\epsilon, 1+\epsilon) \subset A" and "\epsilon>0," we conclude that "1\notin A^\circ". By analogy, "2\notin A^\circ". Therefore, the interior "A^\circ" of "A" is emptyset. Since "\mathbb R\setminus A=(-\infty,1)\cup (1,2)\cup(2,+\infty)" is an open set in usual topology, "A" is a closed set, and thus "\overline A=A". So, "ext(A)=\mathbb R\setminus\overline A=(-\infty,1)\cup (1,2)\cup(2,+\infty)" and the boundary "b(A)=\overline A\setminus A^\circ=A\setminus\emptyset=A."


Since for each "k\in\mathbb Z" there is no basic neigbourhood "(k-\epsilon, k+\epsilon)" such that "(k-\epsilon, k+\epsilon) \subset \mathbb Z" and "\epsilon>0," we conclude that "k\notin \mathbb Z^\circ". Therefore, the interior "\mathbb Z^\circ" of "\mathbb Z" is emptyset. Since "\mathbb R\setminus \mathbb Z=\bigcup_{k\in\mathbb Z}(k,k+1)" is an open set in usual topology, "\mathbb Z" is a closed set, and thus "\overline \mathbb Z=\mathbb Z". So, "ext(\mathbb Z)=\mathbb R\setminus\overline \mathbb Z=\bigcup_{k\in\mathbb Z}(k,k+1)" and the boundary "b(\mathbb Z)=\overline \mathbb Z\setminus \mathbb Z^\circ=\mathbb Z\setminus\emptyset=\mathbb Z."